Find the equation for the line that passes through the point (-1,4) and that is perpendicular to the line with the equation y=1/2x-5/2
Answer by nerdybill(7384)

Find the equation for the line that passes through the point (-1,4) and that is perpendicular to the line with the equation y=1/2x-5/2
.
For a line to be perpendicular the slopes must be "negative reciprocal" of each other.
From the given equation, by inspection, the slope is 1/2.
Our new slope (m) must be:
m(1/2) = -1
m = -2
.
Plug in the slope (-2) and the given point (-1, 4) into the "point-slope" form:
y - y1 = m(x - x1)
y - 4 = -2(x - (-1))
y - 4 = -2(x + 1)
y - 4 = -2x - 2
y = -2x + 2 (this is what they're looking for)